-2y-3(y+1) = -(5y+3)+ y

Answer by yhtong(11)

first, solve all that you can solve,
-2y-3(y+1) = -(5y+3)+ y
-2y-3y-3 =-5y-3+y
-5y-3 =-4y-3
Bring the unknown to the left hand side(LHS), then, the numbers to the right hand side(RHS)
-5y+4y=-3+3
y=0
therefore, y=0